OpenStudy (anonymous):

you roll 2 dice. What is the probability that the total of the two dice is even or that the total is less than 5?

OpenStudy (anonymous):

I get 4/36 or 1/9 - is that correct?

OpenStudy (anonymous):

36 total outcomes minus the odd dice minus the total outcomes that are 5 or higher

OpenStudy (anonymous):

not sure I understand, but I will give you a medal for trying to explain it - thanks!

OpenStudy (anonymous):

1 2 3 4 5 6 1 x x x 2 x x 3 x 4 5 6 Using the graph from earlier, this is the combinations that will give you 5 or less. It's 6 combinations, the other 30 will not give you what the problem requires.

OpenStudy (anonymous):

odd/even=18/36 Less than 5=6/36 Multiplied together it's 1/12. Then again, I have just realized your question is an "OR" question and I'm trying to figure out parameter A AND parameter B.